# Modular Fluctuations in Cosmology

### v1

@@ Abstract
Verlinde and Zurek (VZ) have proposed that quantum gravity fluctuations in causal diamonds lead to observable effects. In particular, they argued that in quantum gravity causal diamonds have an uncertainty in their size that scales as $\delta L\sim\sqrt{\ell_{p}L}$, i.e. fluctuations are enhanced from the Planck scale. In this work, we explore the origin of the VZ scaling by studying scalar perturbations in inflationary cosmology. We reinterpret these cosmological perturbations as fluctuations of the modular Hamiltonian associated with a causal diamond, establishing a connection between modular fluctuations and observable primordial perturbations. Independent of assumptions about the modular Hamiltonian or the background spacetime, our results show that—whenever an (effective) spherically symmetric perturbation is quantized—it leads to the VZ scaling. However, whether such a mode exists in the vacuum of any theory of quantum gravity remains an open question.
@@ 1 Introduction
Our work therefore clarifies how the VZ scaling can come about: by quantizing an (effective) spherically symmetric perturbation in quantum gravity. Because our cosmological setup involves a matter contribution, the s-wave of the inflaton plays this role in inflation. It is unclear to us, however, whether such a mode is also present in the (quantum gravity) vacuum. [^3] In the classical theory, such a mode can be gauged away. This does not preclude its existence in quantum gravity however. If so, our results support the idea that quantum gravity fluctuations are observable at interferometers.
The rest of this paper is organized as follows. In Sec. 2 we review some basic facts about inflationary cosmology and show how primordial scalar perturbations can be related to modular Hamiltonian fluctuations. Independent of assunptions about the modular Hamiltonian, we study the impact of fluctuations on correlators of proper distance and time in Sec. 3. We discuss our results in Sec. 4 and summarize details about the coordinate systems we used in Appendix A and relate different gauges we use to study perturbations in Appendix B.
@@ 2.1 Classical Friedmann-Robertson-Walker Cosmology
The background geometry of the early universe is to good approximation homogeneous and isotropic and described by the Friedmann-Robertson-Walker line element.
@@
where we are working with the conformal time $\eta\in(-\infty,0)$, and $\mathrm{d}\Omega_{2}^{2}$ describes a round two-sphere. Because of the flat spatial slices will refer to this coordinate system as planar coordinates. In planar coordinates, due to the conformal factor the geometry is expanding as the proper distance $\zeta$ between two events with coordinate (or comoving) distance $\rho=\sqrt{\delta_{ij}x^{i}x^{j}}$ scales with time as $\zeta=a(\eta)\rho$.
@@
Because the background geometry only slowly evolves with time it is appropriate to think of each stage of inflation as being described by a de Sitter universe with a slightly different Hubble parameter $H$. Doing so, we can introduce static coordinates in which the metric (2.1) is given by [^4] See Appendix A for an overview of the different coordinates systems we use and their relations.
@@
$B_{i}$ $=\partial_{i}B$ (2.6)
$E_{ij}$ $=C\delta_{ij}+\left(\partial_{i}\partial_{j}-\delta_{ij}\square_{(3)}\right)E~{}.$
Here $\square_{(3)}$ is the Laplacian on flat Euclidean three-space. The metric contains four scalar degrees of freedom $(A,B,C,E)$, but only two of them are physical. Indeed, writing the metric as $g_{\mu\nu}=\bar{g}_{\mu\nu}+h_{\mu\nu}$, with $\bar{g}_{\mu\nu}$ the background metric, we can perform a linear diffeomorphism of the form
@@
$\mathrm{d}s^{2}$ $=b(t,r)^{2}\left(-f(r)\mathrm{d}t^{2}+g(r)^{-1}\mathrm{d}r^{2}+R^{2}\mathrm{d}\Omega_{2}^{2}+\frac{4Hr(\Psi+\Phi)}{1-H^{2}r^{2}}\mathrm{d}t\mathrm{d}r\right)$ (2.10)
@@
$R^{2}$ $=r^{2}(1-2\Phi)~{},$ (2.11)
$f(r)$ $=\left(1-H^{2}r^{2}(1-2\Phi)+2\Psi\right)~{},$
$g(r)$ $=\left(1-H^{2}r^{2}(1-2\Psi)+2\Phi\right)~{}.$
In the absence of anisotropic stress, which is the case for inflation, the equations of motion set $\Psi=\Phi$. Using this, we find that the remaining perturbation $\Phi$ has several effects on the static line element. First, to linear order in $\Phi$ it shifts the root of $f(r)$ to $r_{h}=H^{-1}(1+2\Phi)+{\cal O}(\Phi^{2})$ indicating a shift in the horizon. [^5] We give a more precise definition of the horizon in an FRW geometry in Sec. 2.3. Second, it introduces an off-diagonal term that, in static coordinates [25], indicates a flux of energy through the horizon. In fact, it can be shown that this flux obeys thermodynamic identities that are equivalent to the usual Friedmann equations governing the evolution of the inflationary solution [26].
@@ 2.2 Quantization of Scalar Perturbations
We note that the solution (2.25) is valid for arbitary $k$. By remembering that $u_{k}=z{\cal R}_{k}$ has a straightforward relation with the comoving curvature perturbation and by employing the identity (2.16) in momentum space:
@@ 2.3 Thermodynamics and the Modular Hamiltonian
Now consider what happens when we turn on dynamical gravity. In this case, recent works on the von Neumann (vN) algebra of observables in de Sitter space [32, 33] have revealed that in gravitational systems the type III vN algebra associated to quantum field theory is modified to a type II vN algebra. [^6] To obtain a non-trivial algebra in de Sitter space requires introducing an observer to break the time translation invariance, but in inflationary geometries the inflaton itself already does this [34]. The upshot is that this allows one to rigorously show that the entanglement entropy of some subregion equals the (finite) generalized entropy:
@@
We note that this result reproduces the proportionality of $\langle\Delta K^{2}\rangle\sim\langle K\rangle$ obtained by [5] in anti-de Sitter space and [20] in de Sitter space for a CFT (see [43, 44, 45] for earlier work), but with a prefactor different from one that depends on the renormalization group scale $\mu$. This is not surprising, as the theory we are studying is not conformally invariant. We should also note that the fluctuations of the modular Hamiltonian go to zero in the exact ($\varepsilon\to 0$) de Sitter limit, different from [20]. As explained before, from a cosmologist’s perspective this reflects the fact that scalar perturbations become pure gauge in this limit. However, from the perspective of the modular Hamiltonian this is a consequence of our assumption that the density matrix of pure de Sitter is maximally mixed. We stress that [20] (see also [46]) made a different assumption for the density matrix.
So far we focused on causal diamonds that conisisted of the entire static patch. We are also interested in smaller subregions for which the expectation value of the modular Hamiltonian is still given by the generalized entropy [47]. Using the same identification between the modular Hamiltonian and the Misner-Sharp mass we find that for a smaller subregion with a spherical entangling surface of radius $R$ to leading order in $\varepsilon$
@@ 3.1 Proper Correlators in Cosmology
This simple observation has important consequences. In Sec. 2.3 we saw that in quasi-de Sitter space the correction to the horizon is given by $\delta r=2\Phi$. Treating $\Phi$ as an operator and defining $\delta\zeta=\zeta-\frac{\pi}{2H}$ we see that the first non-zero correlation function is a four-point function [^8] The one-point function $\langle\Phi\rangle$ vanishes.
@@
From this relation we see that the bifurcation surface at $r_{h}=H^{-1}(1+2\Phi)$ is shifted to $(\eta,x)=(\eta_{h},0)$ with
@@
This proper distance can now be easily translated to a proper time by continuing the geometry to the future beyond the cutoff surface and attaching a flat space region. This can be thought off as a (crude) model describing the transition from inflation to a geometry without accelerated expansion and without perturbations. We then consider the same timelike observer following an $x=0$ trajectory that measures the proper time between the null geodesic leaving $x=0$ and returning at $x=0$ in the flat space geometry. To this end, it is useful to write the (unperturbed) planar metric in terms of the time coordinate $\tau=-H^{-1}\log(-H\eta)$ which is the proper time of an observer at $x=0$ (in the unperturbed geometry). In the flat space region, $\tau$ is the standard (proper) Minkowksi time coordinate, see Figure 4.
@@ 3.2 Comments on Other Spacetimes
We recognize this as the static metric up to a conformal factor $C(t,r)$. We thus find that this transformation maps the static patch of de Sitter space to a finite causal diamond with a spherical entangling surface of radius $r=L$ in flat space. We can now consider a linear perturbation to the bifurcation surface in de Sitter space and see how this maps to a change in the bifurcation surface in flat space. Using (3.18) we find that a spherically symmetric perturbation $r=L(1-\delta r)$ of the bifurcation surface at $t=0$ maps to $\tau=0$ and
@@
We see that the leading correction to the location of the bifurcation surface in flat space scales with a square root. Because we are working in flat space, this implies that the proper round-trip time of a photon leaving the observer at $\rho=0$, traveling along the horizon and returning to the observer will have a correction consistent with the VZ scaling. As before, this assumes the presence of a gravitational s-wave perturbation.
Let us contrast this result with the recent work [21]. In that paper, it was demonstrated that quantized graviton fluctuations lead to uncertainty in the arm length of an interferometer that scales as $\delta L^{2}\sim\ell_{p}^{2}$, i.e. the conventional way. From our analysis, we now understand the origin of this different scaling. Employing transverse traceless gauge, [21] focused on quantizing the two physical polarizations of the graviton, i.e. the radiative degrees of freedom in the gravitational field. This leads to a correction to the proper distance of the arm that scales linearly with the perturbation. Instead, inspired by cosmology we quantized a perturbation that modifies the size of the horizon in a spherically symmetric manner. As we demonstrated in (quasi-)de Sitter space, anti-de Sitter space and flat space, this leads to a modification of the proper distance that scales with a square root.
@@
In inflationary cosmology, an s-wave mode is present due to the non-zero matter sector. However, as mentioned before, the presence of such a mode in the vacuum of quantum gravity is less clear. Additionally, while we have shown that correlators of proper distance and time exhibit the VZ scaling, we have not directly related these correlators to an interferometry setup. In particular, it would be worthwhile to better understand the relation between (a change in) the bifurcation surface and the (change in) the armlength of an interferometer. Clearly, this desires future study.
@@ 4 Discussion
Independent of assumptions about modular Hamiltonians, we studied the effect of the spherically symmetric perturbation on proper length and proper time correlators in cosmology. We found that fluctuations in the proper length correlator scale as $\delta\zeta^{2}\sim\ell_{p}/H$. This scaling agrees with [4]. In addition, using a crude model describing the end of inflation in which cosmic expansion and perturbations terminate at a finite time, we demonstrated that proper time fluctuations obey the same scaling, again consistent with the VZ effect.
Furthermore, using the fact that the coordinate system of a finite causal diamond in flat space is conformal to the static de Sitter metric, we provide evidence for the VZ scaling in flat space as well. Our work therefore provides a bridge between inflationary cosmology and quantum gravity effects that are potentially observable in interferometric setups. Importantly, we emphasize that the scaling that we derived is independent of any assumptions about the modular Hamiltonian. Instead, it arises solely from imposing spherical symmetry on the quantized perturbations, in line with the notion that area is quantized in quantum gravity. In cosmology, this spherically symmetric perturbation is sourced by the scalar field driving inflation. However, it is less clear if such modes are present in the vacuum of a quantum gravity theory, for example in flat space.
Our work therefore suggests that s-wave metric fluctuations are at the heart of the VZ effect. A natural and interesting direction for future work is to study mechanisms to generate them. One promising approach is to make use of recent advances in asymptotic symmetries and soft modes. In particular, [19] found an asymptotic charge of a causal diamond in flat space that changes its size. A similar rich story of soft modes and asymptotic symmetries has been studied in cosmology, see e.g. [52], and it would worthwhile to see if there are more lessons about the VZ effect that can be extracted from cosmology.

### v2

@@ Abstract
Verlinde and Zurek (VZ) have proposed that quantum gravity fluctuations in causal diamonds lead to observable effects. In particular, they argued that in quantum gravity causal diamonds have an uncertainty in their size that scales as $\delta L\sim\sqrt{\ell_{p}L}$, i.e. fluctuations are enhanced from the Planck scale. In this work, we explore the origin of the VZ scaling by studying scalar perturbations in inflationary cosmology. We reinterpret these cosmological perturbations as fluctuations of the modular Hamiltonian associated with a causal diamond, establishing a connection between modular fluctuations and observable primordial perturbations. Independent of assumptions about the modular Hamiltonian or the background spacetime, our results show that—whenever an (effective) spherically symmetric perturbation is quantized—it can lead to the VZ scaling. However, whether such a mode exists in the vacuum of any theory of quantum gravity remains an open question.
@@ 1 Introduction
Our work therefore clarifies how the VZ scaling can come about: by quantizing an (effective) spherically symmetric perturbation in quantum gravity. Because our cosmological setup involves a matter contribution, the s-wave of the inflaton plays this role in inflation. It is unclear to us, however, whether such a mode is also present in the (quantum gravity) vacuum. [^3] In the classical theory, such a mode can be gauged away. This does not preclude its existence in quantum gravity however. If so, our results support the idea that quantum gravity fluctuations might be observable at interferometers.
The rest of this paper is organized as follows. In Sec. 2 we review some basic facts about inflationary cosmology and show how primordial scalar perturbations can be related to modular Hamiltonian fluctuations. Independent of assumptions about the modular Hamiltonian, we study the impact of fluctuations on correlators of proper distance and time in Sec. 3. We discuss our results in Sec. 4 and summarize details about the coordinate systems we used in Appendix A and relate different gauges we use to study perturbations in Appendix B.
@@ 2.1 Classical Friedmann-Robertson-Walker Cosmology
The background geometry of the early universe is to a good approximation homogeneous and isotropic and described by the Friedmann-Robertson-Walker line element.
@@
where we are working with the conformal time $\eta\in(-\infty,0)$, and $\mathrm{d}\Omega_{2}^{2}$ describes a round two-sphere. Because of the flat spatial slices, we will refer to this coordinate system as planar coordinates. In planar coordinates, due to the conformal factor the geometry is expanding as the proper distance $\zeta$ between two events with coordinate (or comoving) distance $\rho=\sqrt{\delta_{ij}x^{i}x^{j}}$ scales with time as $\zeta=a(\eta)\rho$.
@@
Because the background geometry only slowly evolves with time it is appropriate to think of each stage of inflation as being described by a de Sitter universe with a slightly different Hubble parameter $H$. Doing so, we can introduce static coordinates in which the metric (2.1) is given by [^4] See Appendix A for an overview of the different coordinate systems we use and their relations.
@@
$B_{i}$ $=\partial_{i}B~{},$ (2.6)
$E_{ij}$ $=C\delta_{ij}+\left(\partial_{i}\partial_{j}-\delta_{ij}\square_{(3)}\right)E~{},$
where $A,\,B,\,C$, and $E$ depend on space and conformal time. Here $\square_{(3)}$ is the Laplacian on flat Euclidean three-space. The metric contains four scalar degrees of freedom $(A,B,C,E)$, but only two of them are physical. Indeed, writing the metric as $g_{\mu\nu}=\bar{g}_{\mu\nu}+h_{\mu\nu}$, with $\bar{g}_{\mu\nu}$ the background metric, we can perform a linear diffeomorphism of the form
@@
$\mathrm{d}s^{2}$ $=b(t,r)^{2}\left(-f\mathrm{d}t^{2}+g^{-1}\mathrm{d}r^{2}+R^{2}\mathrm{d}\Omega_{2}^{2}+\frac{4Hr(\Psi+\Phi)}{1-H^{2}r^{2}}\mathrm{d}t\mathrm{d}r\right)$ (2.10)
@@
$f$ $=\left(1-H^{2}r^{2}(1-2\Phi)+2\Psi\right)~{},$ (2.11)
$g$ $=\left(1-H^{2}r^{2}(1-2\Psi)+2\Phi\right)~{},$
$R^{2}$ $=r^{2}(1-2\Phi)~{}.$
In the absence of anisotropic stress, which is the case for inflation, the equations of motion set $\Psi=\Phi$. Using this, we find that the remaining perturbation $\Phi$ has several effects on the static line element. First, to linear order in $\Phi$ it shifts the root of $f(r)$ indicating a shift in the horizon. [^5] We give the explicit location of the horizon in an FRW geometry in Sec. 2.3. Second, it introduces an off-diagonal term that, in static coordinates [25], indicates a flux of energy through the horizon. In fact, it can be shown that this flux obeys thermodynamic identities that are equivalent to the usual Friedmann equations governing the evolution of the inflationary solution [26].
@@ 2.2 Quantization of Scalar Perturbations
We note that the solution (2.25) is valid for arbitrary $k$. By remembering that $u_{k}=z{\cal R}_{k}$ has a straightforward relation with the comoving curvature perturbation and by employing the identity (2.16) in momentum space:
@@ 2.3 Thermodynamics and the Modular Hamiltonian
Now consider what happens when we turn on dynamical gravity. In this case, recent works on the von Neumann (vN) algebra of observables in de Sitter space [32, 33] have revealed that in gravitational systems the type III vN algebra associated to quantum field theory is modified to a type II vN algebra. [^6] To obtain a non-trivial algebra in de Sitter space requires introducing an observer to break the time translation invariance, but in inflationary geometries the inflaton itself already does this [34]. Related recent work [35] further explores the role of spacetime translation symmetry breaking in the construction of gauge-invariant operators. The upshot is that this allows one to rigorously show that the entanglement entropy of some subregion equals the (finite) generalized entropy:
@@
We note that this result reproduces the proportionality of $\langle\Delta K^{2}\rangle\sim\langle K\rangle$ obtained by [5] in anti-de Sitter space and [20] in de Sitter space for a CFT (see [44, 45, 46] for earlier work), but with a prefactor different from one that depends on the renormalization group scale $\mu$. This is not surprising, as the theory we are studying is not conformally invariant. We should also note that the fluctuations of the modular Hamiltonian go to zero in the exact ($\varepsilon\to 0$) de Sitter limit, different from [20]. As explained before, from a cosmologist’s perspective this reflects the fact that the inflaton fluctuations become pure gauge in this limit. However, from the perspective of the modular Hamiltonian this is a consequence of our assumption that the density matrix of pure de Sitter is maximally mixed. We stress that [20] (see also [47]) made a different assumption for the density matrix.
So far we focused on causal diamonds that consisted of the entire static patch. We are also interested in smaller subregions for which the expectation value of the modular Hamiltonian is still given by the generalized entropy [48]. Using the same identification between the modular Hamiltonian and the Misner-Sharp mass we find that for a smaller subregion with a spherical entangling surface of radius $R$ to leading order in $\varepsilon$
@@ 3.1 Proper Correlators in Cosmology
This simple observation has important consequences. In Sec. 2.3 we saw that in quasi-de Sitter space the radius of the horizon reads $R_{h}=H^{-1}(1+\Phi)$, which corresponds to $\rho_{h}=-\eta(1+2\Phi)$ in terms of the coordinate distance in planar coordinates. [^8] Note that the relation between the radius of a sphere $R$ and the coordinate distance $\rho$ in perturbed quasi-de Sitter space is given by $R^{2}=a(\eta)^{2}\rho^{2}(1-2\Phi)$. By using the coordinate transformation (A.7), the correction to the horizon in static coordinates is given by $\delta r=2\Phi$. Treating $\Phi$ as an operator and defining $\delta\zeta=\zeta-\frac{\pi}{2H}$ we see that the first non-zero correlation function is a four-point function [^9] The one-point function $\langle\Phi\rangle$ vanishes.
@@
From this relation we see that the bifurcation surface at $r_{h}=H^{-1}(1+2\Phi)$ is shifted to $(\eta,x)=(\eta_{h},0)$ with [^12] It is important to note that we consider perturbations of the form (2.10) in Milne-like coordinates. Defining the bifurcation surface using the (conformal) Killing vector $\partial_{t}$ this leads to a change in the bifurcation surface of $r_{h}=H^{-1}(1+2\Phi)$.
@@
This proper distance can now be easily translated to a proper time by continuing the geometry to the future beyond the cutoff surface and attaching a flat space region. This can be thought off as a (crude) model describing the transition from inflation to a geometry without accelerated expansion and without perturbations. We then consider the same timelike observer following an $x=0$ trajectory that measures the proper time between the null geodesic leaving $x=0$ and returning at $x=0$ in the flat space geometry. To this end, it is useful to write the (unperturbed) planar metric in terms of the time coordinate $\tau=-H^{-1}\log(-H\eta)$ which is the proper time of an observer at $x=0$ (in the unperturbed geometry). In the flat space region, $\tau$ is the standard (proper) Minkowski time coordinate, see Figure 4.
@@ 3.2 Comments on Other Spacetimes
We recognize this as the static metric up to a conformal factor $C(t,r)$. We thus find that this transformation maps the static patch of de Sitter space to a finite causal diamond with a spherical entangling surface of radius $r=L$ in flat space. We can now consider a linear perturbation to the bifurcation surface in de Sitter space and see how this maps to a change in the bifurcation surface in flat space.
Using (3.18) we find that a spherically symmetric perturbation $r=L(1-\delta r)$ of the bifurcation surface at $t=0$ maps to $\tau=0$ and
@@
We see that the leading correction to the location of the bifurcation surface in flat space scales with the square root of the perturbation. Because we are working in flat space, this implies that the proper round-trip time of a photon leaving the observer at $\rho=0$, traveling along the horizon and returning to the observer also has a correction that scales with the square root of the perturbation consistent with the VZ scaling.
As before, this assumes the presence of a gravitational s-wave perturbation that shifts the bifurcation surface of conformal-de Sitter space linearly. We’d like to stress that this assumption involves a choice of conformal Killing vector that specifies the perturbed causal diamond. It is possible, for example, to consider a conformal Killing vector for which the coordinate location of the bifurcation surface doesn’t change in the perturbed metric.
Let us contrast the square root scaling of the round-trip time with the recent work [21]. In that paper, it was demonstrated that quantized graviton fluctuations lead to uncertainty in the arm length of an interferometer that scales as $\delta L^{2}\sim\ell_{p}^{2}$, i.e. the conventional way. From our analysis, we now understand the origin of this different scaling. Employing transverse traceless gauge, [21] focused on quantizing the two physical polarizations of the graviton, i.e. the radiative degrees of freedom in the gravitational field. This leads to a correction to the proper distance of the arm that scales linearly with the perturbation. Instead, inspired by cosmology we quantized a perturbation that modifies the size of the horizon in a spherically symmetric manner. As we demonstrated in (quasi-)de Sitter space, anti-de Sitter space and flat space, this leads to a modification of the proper distance that scales with a square root.
@@
In inflationary cosmology, an s-wave mode is present due to the non-zero matter sector. However, as mentioned before, the presence of such a mode in the vacuum of quantum gravity is less clear. Additionally, while we have shown that correlators of proper distance and time exhibit the VZ scaling, we have not directly related these correlators to an interferometry setup. In particular, it would be worthwhile to better understand the relation between (a change in) the bifurcation surface and the (change in) the armlength of an interferometer. Clearly, this deserves future study.
@@ 4 Discussion
Independent of assumptions about modular Hamiltonians, we studied the effect of the spherically symmetric perturbation on proper length and proper time correlators in cosmology. We found that fluctuations in the proper length correlator scale as $\delta\zeta^{2}\sim\ell_{p}/H$. This scaling agrees with [4]. Using a crude model describing the end of inflation in which cosmic expansion and perturbations terminate at a finite time, we demonstrated that proper time fluctuations obey the same scaling, again consistent with the VZ effect.
Interestingly, the geometric mean of the horizon radius and the Planck scale has also appeared as the scale relevant to certain quantum gravity effects in [53, 54, 55, 1, 56, 57]. The origin of this scaling seems to lie in a fundamental limitation on the number of degrees of freedom in a subregion given by the Bekenstein bound [58] or its covariant version (the Bousso bound [59, 60]). In particular, in [54] the geometric mean of the Planck scale and inverse Hubble parameter arose due to a bound on the degrees of freedom in phase space, whereas in our work we directly studied fluctuations in the area (proportional to the entropy) of a cosmological causal diamond. To better understand the relation between the VZ effect and the setup of [54], it would be interesting to examine the form of the density matrix used in [54]. This is a necessary step, because just satisfying a covariant entropy bound is insufficient to determine the precise form of the density matrix [20]. Knowing the density matrix, it is then possible to compute the associated modular Hamiltonian and see if its fluctuations scale as $\langle\Delta K^{2}\rangle\sim\langle K\rangle$, as expected for the VZ effect.
Furthermore, using the fact that the coordinate system of a finite causal diamond in flat space is conformal to the static de Sitter metric, we provide evidence for the VZ scaling in flat space as well, assuming a perturbation that shifts the bifurcation surface of conformal-de Sitter space linearly. Our work therefore provides a bridge between inflationary cosmology and quantum gravity effects that are potentially observable in interferometric setups. Importantly, we emphasize that the scaling that we derived is independent of any assumptions about the modular Hamiltonian. Instead, it arises solely from imposing spherical symmetry on the quantized perturbations, in line with the notion that area is quantized in quantum gravity. In cosmology, this spherically symmetric perturbation is sourced by the scalar field driving inflation. However, it is less clear if such modes are present in the vacuum of a quantum gravity theory, for example in flat space.
Our work therefore suggests that s-wave metric fluctuations are at the heart of the VZ effect. A natural and interesting direction for future work is to study mechanisms to generate them. One promising approach is to make use of recent advances in asymptotic symmetries and soft modes. In particular, [19] found an asymptotic charge of a causal diamond in flat space that changes its size. A similar rich story of soft modes and asymptotic symmetries has been studied in cosmology, see e.g. [61], and it would be worthwhile to see if there are more lessons about the VZ effect that can be extracted from cosmology.